OpenStudy (anonymous):

3x + 2 = y; y = 2 whats x??

OpenStudy (anonymous):

Substitute y = 2 in 3x + 2 = y we find: 3x + 2 = 2 i.e. 3x= 2-2 i.e. 3x= 0 i.e. x=0/3 =?

OpenStudy (anonymous):

@Rico050

OpenStudy (anonymous):

ohhhh!!! so x equals zero!

OpenStudy (anonymous):

Yes exactly...

OpenStudy (anonymous):

x=0